Shift summary for owl shift, 8/Jan/00


Subject: Shift summary for owl shift, 8/Jan/00
From: Vivian O'Dell (odell@fnal.gov)
Date: Fri Feb 04 2000 - 15:10:18 CST


               Shift summary for owl shift, 8/Jan/00

On Shift: MT, AG

Hours of beam available at all intensities: 6.8h
Hours of beam available at nominal intensity: 6.8h
Hours of E799 data (special runs not included): 5.9h
Hours of special runs (laser scans, muon runs...): 0.0h

Accounting of lost time:

o twice got screwed up L3 accept fraction; each time run was stopped and
  dart was fresh-started
o minor accelerator glitches.

o tape change and problems starting new run afterwards.

========================================================================

Runs ended this shift:

15433 - 7.3M - BEAM_799
15434 - 3.3M - BEAM_799

Tapes brought to FCC this shift:
  (including tapes for on-line split)

none.

========================================================================

Accesses during shift:

none.

========================================================================

Overview of shift:

Pretty smooth data-taking with two short accelerator glitches and about
half an hour of beam lost due to L3 accept fraction disorder, which
happened twice during this shift. Also lost last 25 minutes of the shift
due to tape change and problems while starting a new run.

========================================================================

Description of Shift:

00:24 Sydney did Hard Reboot of krpipe11 to get rid of "Message
                from Pipe-line" message, which was triggered every 11 min
                40 sec.

00:24 Just at this time vacuum in Booster bursts.

00:30 spill 113 Beam's back.

00:50 Mounted UPF074 on ktev4.

01:30 spill 157 L3 accept fraction problem. Just before that HCC was
                inefficient for couple of spills.
                -> stopped the run, updated ADGN as per request on the
                white board.

01:46 Run 15434 started.

03:10 spill 61 Whole CsI quadrant (upper left on the laser monitoring
                histogram) is missing; also on event display upper right
                quadrant has no data or cosmics. The quadrant is missing
                for a few spills and L3 accept fraction is screwed up
                second time for this shift. Paged Julie and she suggested
                resetting pipe-line and fresh-starting if that doesn't
                help.
                -> end the run

03:30 Started run 15435.

03:35 Onsplit tape UPH055 is mounted on ktev2.

04:40 Onsplit disks /ktev1,2,4g are 91% full ...

05:15 ... 92% full

05:16 spill 81 Beam is gone due to Klystron 1 RF problem.

05:22 spill 86 Beam is back.

05:40 /ktev1,2,4g disks are 93% full, so onsplit definitely
                cannot cope with it -> page Arun (although he left
                control room around 1am... ). He'll delete some files
                from home.

06:35 Tapes are full. Changed tapes, but run 15436 failed to
                start (after successful initialization).

06:45 Dart fresh start, but now we even cannot pass the
                initialization. After some problems with tape drive
                and help from Ed the run was finally started (details are
                coming in the day shift summary).

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift



This archive was generated by hypermail 2b27 : Fri Feb 04 2000 - 15:10:21 CST